OpenStudy (anonymous):

Another question--similar. I did try, but couldn't get it right for some reason.

OpenStudy (anonymous):

what did you do ?

hartnn (hartnn):

what is 11/33 = ?

hartnn (hartnn):

\(\huge \frac{a^m}{a^n}=a^{m-n}\)

hartnn (hartnn):

keep it as 1/3

hartnn (hartnn):

\(\huge \frac{d^2}{d^4}=d^{2-4}\) when division, you subtract the exponents

hartnn (hartnn):

then what will be e^5/e^2 ?

hartnn (hartnn):

that is correct!

hartnn (hartnn):

e^5/e^2 = e^{5-2} = e^3

hartnn (hartnn):

and the exponent of d u got ?

hartnn (hartnn):

yup, but thats in denominator. now there is ^2 overall, so just double all the exponents, what u get?

hartnn (hartnn):

yes double. 3*2 = 6 and not 9 so u get e^6 d^4 and whats the square of (1/3) ?

hartnn (hartnn):

no, square (1/3)(1/3)=?

hartnn (hartnn):

yup, 9 in denominator. so which option is it ?

hartnn (hartnn):

yup

hartnn (hartnn):

sure

OpenStudy (anonymous):

it becomes \[\frac{ (-4)^3(x^2)^3 }{ x^3(z^2)^3 }\]

OpenStudy (anonymous):

what will you get then ?

OpenStudy (anonymous):

first you forgot z.. the second thing is that at the numerator you got (x^2)^3 which is x^(2*3) in the denominator you got only x^3

OpenStudy (anonymous):

so you meant to write on the denominator z^6 instead of x^6 .. you just wrote it wrong but if thats what you meant so it is correct \[\frac{ -64x^3 }{ z^6 }\]

OpenStudy (anonymous):

:)